# Efficient and Scalable Monocular Human-Object Interaction Motion Reconstruction

### v1

@@ 1 Introduction
Despite its importance, acquiring generalized 4D HOI data remains a significant challenge. Existing multi-sensor oriented HOI capture systems, exemplified by BEHAVE [2], offer high-precision reconstructions by leveraging multi-view camera setups and sophisticated sensor arrays. However, the associated cost, from numerous cameras, sensors, and specialized studios, is prohibitive for large-scale data collection. Consequently, these systems are often confined to controlled indoor settings with limited object diversity, preventing the capture of HOI data for common outdoor activities like riding and surfing, as well as complex on-site construction tasks involved in industrial manufacturing processes.
Recognizing the limitations of high-precision capture systems, recent research has explored reconstructing 3D HOI from widely available monocular images. For example, Open3DHOI [35] relies on manual object pose adjustment in Blender, while PICO [6] proposes annotating contact regions. However, while effective for static images, both annotation strategies become prohibitively expensive and time-consuming when applied to video sequences, and they fundamentally struggle to enforce spatio-temporal consistency across frames. To address these critical challenges, we propose 4DHOISolver, a novel and efficient framework for generating high-quality, temporally-consistent 4D HOI data from diverse monocular internet videos.
@@ 2 Related Works
Recently, many works [10, 2, 35, 6, 25, 33, 43, 30, 26, 31, 27] have focused on constructing 3D and 4D HOI datasets. Some works use multi-view RGB-D cameras to reconstruct human SMPL [18] / SMPL-X [21] motion and object poses from depth point clouds and Mocap. BEHAVE [2] set up and calibrated 4 RGB-D cameras at four corners in a square recording volume, involving 8 subjects. InterCAP [25] followed BEHAVE by adding hand reconstruction and 2 cameras. Other works, such as HODome [27], used 76 RGB cameras for data capture, and $HOI-M^{3}$ [33] used 42 cameras. Other methods use Mocap suits for data collection, such as OMOMO [30], HUMOTO [26], and TRUMANS [31]. For object tracking, BEHAVE used fixed object keypoints to annotate positions in multi-view images manually, OMOMO and TRUMANS used markers, HUMOTO used two cameras, and FoundationPose [34] for tracking.
With the advancement of 3D reconstruction techniques, some recent works  [43, 35, 6, 41, 9] have started to explore reconstructing human-object interaction data from monocular images and videos. Most of these works obtain human motion using state-of-the-art motion reconstruction tools, while object motion is handled with a variety of different approaches. WildHOI [43] constrains object pose by using pre-defined keypoints on object templates and adjusting their positions. Open3DHOI [35] reconstructs object meshes using state-of-the-art 3D reconstruction tools, followed by refinement in Blender. PICO [6] determines the object by matching the target image with similar meshes from a 3D object dataset [8] and applies contact constraints for further refinement.
@@
3D reconstruction tools provide foundations for reconstructing human-object interactions from monocular videos. First, human pose reconstruction tools [16, 7, 13, 24, 28] have become quite mature and can accurately estimate human poses in most scenarios. OSX [16] can simultaneously predict the body pose, hand pose, and facial expression of a person, but it lacks global information. GVHMR [7] can effectively reconstruct human motion in world space while ensuring alignment with the in-camera coordinate system and the input video, making it highly suitable for our project; however, it is unable to predict hand poses. HAMER [28] specializes in hand pose estimation, capable of accurately capturing hand details, which facilitates its integration with GVHMR for full-body pose estimation.
Secondly, 3D generative models [36, 23, 37] have made significant progress, especially in reconstructing target objects from single or multiple images, enabling the generation of high-quality meshes. These models are trained on large-scale 3D datasets like Objaverse [8], which can perform well on real images with high resolution now. In addition, image inpainting techniques [20, 4] can help address cases with partial occlusions. Pix2Gestalt [20] uses diffusion to predict the occluded parts of objects from images, enabling simultaneous prediction of both the masked regions and the corresponding RGB information. Diffusion-VAS [4] can complete occluded parts of objects in videos while maintaining temporal consistency in the reconstructed regions.
@@ 3.1 Data Collection
1) Tracking both human and object masks using SAM2 [15]. For TikTok data, we also need to detect shot transitions and select the first frame containing human–object interaction as the starting point for subsequent reconstruction.
2) Next, we use image inpainting method [20] to complete the occluded object regions. Then we use TRELLIS [36] for obejct reconstruction.
3) For human motion reconstruction, we employ GVHMR [7] for the full-body SMPL-X pose and HAMER [28] for fine-grained hand motion. The fused poses can ensure the plausibility of motion in most data.
4) Initializing and aligning the HOI poses through a depth-aware projection approach. Specifically, we estimate the depth of humans and objects based on DepthAnythingv2 [40] and use the depth point cloud for the spatial alignment and object scale estimation by projecting the human mesh and object mesh together to fit the point cloud.
@@ 3.3.2 Annotation Procedure
Based on the fact that there are usually some fixed points in HOI motion, we aim to find and track these points, which makes the annotation process easy and fast. The user only needs to re-annotate when the stable points are changed in the video. Specifically, we use the Point Tracking model [14] to track the annotated 2D points. Interactive objects are divided into movable and static categories. For objects that remain motionless throughout the video, we provide an additional annotation option, allowing users to indicate whether the object pose should be fixed based on the video content. What’s more, for cases in Sec. 3.2 where the reconstructed object scale is inaccurate, we adjust the object’s scale before annotation.
@@ 3.4.2 HOI Optimizer
$L$ address the limitation of HOI Keypoint Solver [11]’s point-based optimization, which ignores the physical plausibility of interaction, such as penetration, and allows for more fine-grained adjustments of the object’s pose estimation.
@@ 4 HOI Simulation
To demonstrate the scalability and easy-to-use property of our data, we adopt a test via humanoid body control. We train both the humanoid and the object in simulation with reinforcement learning. However, existing methods [19, 32, 29, 39] usually impose stringent requirements on the quality of the reference motion. As our reference motion is reconstructed from monocular videos with limited constraints, unreasonable reference HOI poses are sometimes inevitable under unseen or missing viewpoints, which affects the simulation results of current methods.
To this end, we developed a noise-robust, reference-less-reliable HOI simulation method based on InterMimic [39]. The goal of HOI simulation is to train a policy $\pi$ that generates a physically plausible HOI motion sequence $\{q_{t}^{h},q_{t}^{o}\}_{t=1}^{T}$. This sequence should be close to a reference motion sequence $\{\hat{q}_{t}^{h},\hat{q}_{t}^{o}\}_{t=1}^{T}$, which is reconstructed by our 4DHOISolver.
@@ 5.1.1 Diversity
Interaction Diversity. We used the Qwen2.5-vl-72B [1] model to annotate actions in the videos and further manually verified them. Ultimately, our dataset contains 103 object categories. For the BEHAVE and OMOMO datasets, we used Qwen2.5-72B [22] to extract action categories from the textual descriptions in the dataset and manually filtered the results. The outcomes are summarized in Tab. 2.
@@ 5.1.2 Data Quality
Reconstruction Quality. We evaluated the quality of object reconstruction by comparing the similarity between the reconstructed objects and the images as the metric. We used OpenShape[17] to compare the feature similarity between meshes and their corresponding images, and the results in Tab. 2 show that our dataset outperforms other open-vocabulary datasets.
@@ 5.2 HOI Motion Reconstruction
The evaluated baselines include our automated reconstruction results (Recon), using FoundationPose for object pose estimation, and InteractVLM [9] with reconstruction performed every 10 frames. We also compared our method without the HOI Optimizer refinement. The results in Tab. LABEL:tab:phys_metrics demonstrate that our full method achieves the best optimization performance. Fig. 7 shows the visualization of our optimization results compared with FoundationPose.
@@ 6 HOI Contact Prediction
We test models including LEMON [41] and InteractVLM [9], which simultaneously predict human and object interaction regions; the 3D point cloud large model PointLLM [38]; and the LLM-based object affordance prediction model 3D-AffordanceLLM [5]. To reduce task difficulty, we use only the recall rate as the evaluation metric, measuring how much of the annotated contact region is correctly covered by each model’s predicted contact region. For 3D Q&A models like PointLLM, we input the merged point cloud of humans and objects optimized by 4DHOISolver, along with the object category, prompting the model to generate a textual description of the focused interaction human regions. We then use Qwen2.5-72B to extract the joint names from this description. The results in Tab. 4 show that existing 3D models still perform poorly in interaction prediction, especially when simultaneously predicting the correspondence between human and object interactions.
@@ A.3.1 TikTok Reconstruction
We follow PICO’s object selection method by using OpenShape [17] from Objaverse to choose the most similar object, then compare its similarity score with the reconstructed object, selecting the higher one as the final object. This helps save bad reconstructions with TRELLIS due to severe occlusion.
@@ A.6.1 Loss
Collision Loss. To prevent interpenetration between the human mesh and the object mesh, we adopt the bidirectional mesh-to-mesh collision penalty used in [12]. Specifically, we apply the same collision operator to measure (i) human vertices inside the object surface and (ii) object vertices inside the human surface. The final collision loss is a weighted combination
@@ B.2 Action Annotation
In Sec. 5.1.1, we mention that we annotated 103 action categories for our dataset. We automatically extracted these action categories using a two-step strategy involving Qwen2.5-VL-72B [1] and Qwen2.5-72B-Instruct [22], followed by manual filtering.
@@
Next, we used Qwen2.5-72B-Instruct to extract the action categories. We adopted the action list from HICO-DE [3] as a pool of candidate actions and prompted the model to select the actions present in each video from this list. The prompt we provided was: “Here is an human-{obj_name} action description: {interaction_description} Please select the most similar human-{obj_name} interaction action from the following candidate list.”
@@ C.1 Implementation Details
The reference data is derived from human bodies represented using SMPL-X [21]. For simulation, we retarget these models into rigid bodies following [19, 42], and Objects are also converted into simulation models through convex decomposition. Similar to existing methods [29, 39], we perform HOI-simulation in Isaac Gym and use the first reference frame to initialize the simulation environment. The parameter settings for simulation environments are shown in the Tab. 7 To distinguish between static and dynamic objects derived from the annotation

### v2

@@ 1 Introduction
Despite its importance, acquiring generalized 4D HOI data remains a significant challenge. Existing multi-sensor oriented HOI capture systems, exemplified by BEHAVE [behave], offer high-precision reconstructions by leveraging multi-view camera setups and sophisticated sensor arrays. However, the associated cost, from numerous cameras, sensors, and specialized studios, is prohibitive for large-scale data collection. Consequently, these systems are often confined to controlled indoor settings with limited object diversity, preventing the capture of HOI data for common outdoor activities like riding and surfing, as well as complex on-site construction tasks involved in industrial manufacturing processes.
Recognizing the limitations of high-precision capture systems, recent research has explored reconstructing 3D HOI from widely available monocular images. For example, Open3DHOI [open3dhoi] relies on manual object pose adjustment in Blender, while PICO [pico] proposes annotating contact regions. However, while effective for static images, both annotation strategies become prohibitively expensive and time-consuming when applied to video sequences, and they fundamentally struggle to enforce spatio-temporal consistency across frames. To address these critical challenges, we propose 4DHOISolver, a novel and efficient framework for generating high-quality, temporally-consistent 4D HOI data from diverse monocular internet videos.
@@ 2 Related Works
Recently, many works [PROX, behave, open3dhoi, pico, intercap, hoim3, wildhoi, omomo, humoto, trumans, hodome] have focused on constructing 3D and 4D HOI datasets. Some works use multi-view RGB-D cameras to reconstruct human SMPL [smpl] / SMPL-X [smpl-x] motion and object poses from depth point clouds and Mocap. BEHAVE [behave] set up and calibrated 4 RGB-D cameras at four corners in a square recording volume, involving 8 subjects. InterCAP [intercap] followed BEHAVE by adding hand reconstruction and 2 cameras. Other works, such as HODome [hodome], used 76 RGB cameras for data capture, and $HOI-M^{3}$ [hoim3] used 42 cameras. Other methods use Mocap suits for data collection, such as OMOMO [omomo], HUMOTO [humoto], and TRUMANS [trumans]. For object tracking, BEHAVE used fixed object keypoints to annotate positions in multi-view images manually, OMOMO and TRUMANS used markers, HUMOTO used two cameras, and FoundationPose [foundationpose] for tracking.
With the advancement of 3D reconstruction techniques, some recent works  [wildhoi, open3dhoi, pico, lemon, interactvlm] have started to explore reconstructing human-object interaction data from monocular images and videos. Most of these works obtain human motion using state-of-the-art motion reconstruction tools, while object motion is handled with a variety of different approaches. WildHOI [wildhoi] constrains object pose by using pre-defined keypoints on object templates and adjusting their positions. Open3DHOI [open3dhoi] reconstructs object meshes using state-of-the-art 3D reconstruction tools, followed by refinement in Blender. PICO [pico] determines the object by matching the target image with similar meshes from a 3D object dataset [objaverseXL] and applies contact constraints for further refinement.
@@
3D reconstruction tools provide foundations for reconstructing human-object interactions from monocular videos. First, human pose reconstruction tools [osx, gvhmr, hmr, aios, hamer] have become quite mature and can accurately estimate human poses in most scenarios. OSX [osx] can simultaneously predict the body pose, hand pose, and facial expression of a person, but it lacks global information. GVHMR [gvhmr] can effectively reconstruct human motion in world space while ensuring alignment with the in-camera coordinate system and the input video, making it highly suitable for our project; however, it is unable to predict hand poses. HAMER [hamer] specializes in hand pose estimation, capable of accurately capturing hand details, which facilitates its integration with GVHMR for full-body pose estimation.
Secondly, 3D generative models [trellis, shi2023zero123plus, xu2024instantmesh] have made significant progress, especially in reconstructing target objects from single or multiple images, enabling the generation of high-quality meshes. These models are trained on large-scale 3D datasets like Objaverse [objaverseXL], which can perform well on real images with high resolution now. In addition, image inpainting techniques [pix2gestalt, diffusionvas] can help address cases with partial occlusions. Pix2Gestalt [pix2gestalt] uses diffusion to predict the occluded parts of objects from images, enabling simultaneous prediction of both the masked regions and the corresponding RGB information. Diffusion-VAS [diffusionvas] can complete occluded parts of objects in videos while maintaining temporal consistency in the reconstructed regions.
@@ 3.1 Data Collection
1) Tracking both human and object masks using SAM2 [sam2]. For TikTok data, we also need to detect shot transitions and select the first frame containing human–object interaction as the starting point for subsequent reconstruction.
2) Next, we use image inpainting method [pix2gestalt] to complete the occluded object regions. Then we use TRELLIS [trellis] for obejct reconstruction.
3) For human motion reconstruction, we employ GVHMR [gvhmr] for the full-body SMPL-X pose and HAMER [hamer] for fine-grained hand motion. The fused poses can ensure the plausibility of motion in most data.
4) Initializing and aligning the HOI poses through a depth-aware projection approach. Specifically, we estimate the depth of humans and objects based on DepthAnythingv2 [depth_anything_v2] and use the depth point cloud for the spatial alignment and object scale estimation by projecting the human mesh and object mesh together to fit the point cloud.
@@ 3.3.2 Annotation Procedure
Based on the fact that there are usually some fixed points in HOI motion, we aim to find and track these points, which makes the annotation process easy and fast. The user only needs to re-annotate when the stable points are changed in the video. Specifically, we use the Point Tracking model [cotracker] to track the annotated 2D points. Interactive objects are divided into movable and static categories. For objects that remain motionless throughout the video, we provide an additional annotation option, allowing users to indicate whether the object pose should be fixed based on the video content. What’s more, for cases in Sec. 3.2 where the reconstructed object scale is inaccurate, we adjust the object’s scale before annotation.
@@ 3.4.2 HOI Optimizer
$L$ address the limitation of HOI Keypoint Solver [hoikeypointsolver]’s point-based optimization, which ignores the physical plausibility of interaction, such as penetration, and allows for more fine-grained adjustments of the object’s pose estimation.
@@ 4 HOI Simulation
To demonstrate the scalability and easy-to-use property of our data, we adopt a test via humanoid body control. We train both the humanoid and the object in simulation with reinforcement learning. However, existing methods [Luo2023PerpetualHC, Wang_2025_CVPR, wang2023physhoiphysicsbasedimitationdynamic, xu2025intermimic] usually impose stringent requirements on the quality of the reference motion. As our reference motion is reconstructed from monocular videos with limited constraints, unreasonable reference HOI poses are sometimes inevitable under unseen or missing viewpoints, which affects the simulation results of current methods.
To this end, we developed a noise-robust, reference-less-reliable HOI simulation method based on InterMimic [xu2025intermimic]. The goal of HOI simulation is to train a policy $\pi$ that generates a physically plausible HOI motion sequence $\{q_{t}^{h},q_{t}^{o}\}_{t=1}^{T}$. This sequence should be close to a reference motion sequence $\{\hat{q}_{t}^{h},\hat{q}_{t}^{o}\}_{t=1}^{T}$, which is reconstructed by our 4DHOISolver.
@@ 5.1.1 Diversity
Interaction Diversity. We used the Qwen2.5-vl-72B [Qwen2.5-VL] model to annotate actions in the videos and further manually verified them. Ultimately, our dataset contains 103 object categories. For the BEHAVE and OMOMO datasets, we used Qwen2.5-72B [qwen2.5] to extract action categories from the textual descriptions in the dataset and manually filtered the results. The outcomes are summarized in Tab. 2.
@@ 5.1.2 Data Quality
Reconstruction Quality. We evaluated the quality of object reconstruction by comparing the similarity between the reconstructed objects and the images as the metric. We used OpenShape[openshape] to compare the feature similarity between meshes and their corresponding images, and the results in Tab. 2 show that our dataset outperforms other open-vocabulary datasets.
@@ 5.2 HOI Motion Reconstruction
The evaluated baselines include our automated reconstruction results (Recon), using FoundationPose for object pose estimation, and InteractVLM [interactvlm] with reconstruction performed every 10 frames. We also compared our method without the HOI Optimizer refinement. The results in Tab. LABEL:tab:phys_metrics demonstrate that our full method achieves the best optimization performance. Fig. 7 shows the visualization of our optimization results compared with FoundationPose.
@@ 6 HOI Contact Prediction
We test models including LEMON [lemon] and InteractVLM [interactvlm], which simultaneously predict human and object interaction regions; the 3D point cloud large model PointLLM [pointllm]; and the LLM-based object affordance prediction model 3D-AffordanceLLM [3d_affordancellm]. To reduce task difficulty, we use only the recall rate as the evaluation metric, measuring how much of the annotated contact region is correctly covered by each model’s predicted contact region. For 3D Q&A models like PointLLM, we input the merged point cloud of humans and objects optimized by 4DHOISolver, along with the object category, prompting the model to generate a textual description of the focused interaction human regions. We then use Qwen2.5-72B to extract the joint names from this description. The results in Tab. 4 show that existing 3D models still perform poorly in interaction prediction, especially when simultaneously predicting the correspondence between human and object interactions.
@@ A.3.1 TikTok Reconstruction
We follow PICO’s object selection method by using OpenShape [openshape] from Objaverse to choose the most similar object, then compare its similarity score with the reconstructed object, selecting the higher one as the final object. This helps save bad reconstructions with TRELLIS due to severe occlusion.
@@ A.6.1 Loss
Collision Loss. To prevent interpenetration between the human mesh and the object mesh, we adopt the bidirectional mesh-to-mesh collision penalty used in [collision]. Specifically, we apply the same collision operator to measure (i) human vertices inside the object surface and (ii) object vertices inside the human surface. The final collision loss is a weighted combination
@@ B.2 Action Annotation
In Sec. 5.1.1, we mention that we annotated 103 action categories for our dataset. We automatically extracted these action categories using a two-step strategy involving Qwen2.5-VL-72B [Qwen2.5-VL] and Qwen2.5-72B-Instruct [qwen2.5], followed by manual filtering.
@@
Next, we used Qwen2.5-72B-Instruct to extract the action categories. We adopted the action list from HICO-DE [hico-det] as a pool of candidate actions and prompted the model to select the actions present in each video from this list. The prompt we provided was: “Here is an human-{obj_name} action description: {interaction_description} Please select the most similar human-{obj_name} interaction action from the following candidate list.”
@@ C.1 Implementation Details
The reference data is derived from human bodies represented using SMPL-X [smpl-x]. For simulation, we retarget these models into rigid bodies following [Luo2023PerpetualHC, yuan2021simpoesimulatedcharactercontrol], and Objects are also converted into simulation models through convex decomposition. Similar to existing methods [wang2023physhoiphysicsbasedimitationdynamic, xu2025intermimic], we perform HOI-simulation in Isaac Gym and use the first reference frame to initialize the simulation environment. The parameter settings for simulation environments are shown in the Tab. 7 To distinguish between static and dynamic objects derived from the annotation